OpenStudy (anonymous):

find the GCF and then factor out the GCF 1.15y-3 2.4x^2+3x 3.6x^3y+7x^2+x

OpenStudy (anonymous):

15y-3=3(5y-1) 4x^2+3x=x(4x+3) 6x^3y+7x^2+x=x(6x^2y+7x+1) GCF is 1